OpenStudy (madhatter):

Write a system of equations for this situation. Currently, John has a collection of 32 CDs and William has a collection of 24 CDs. John is adding 1 CDs per month to his collection, while William plans to add 2 CDs per month. is it a or is it b A.y = x + 32 y = 2x + 24 B.y = x + 24 y = 2x + 32

OpenStudy (anonymous):

well what do you think you do first?

OpenStudy (madhatter):

what do u mean..?

OpenStudy (anonymous):

what do you think one of your equations will be? also you omited the actual question and only gave the background info I think

OpenStudy (madhatter):

no i think i posted what i ment to post and i ment to post a question that has two awnesers to pick from simple as that

jimthompson5910 (jim_thompson5910):

Let's focus on John If x = # of months and y = # of CDs then y = 1x + 32 How did I get this? Well this is in y = mx+b form where m = 1 and b = 32 The slope is 1 because this is how much his CD collection is changing: he's adding 1 CD a month, so that's why the slope is 1 The y-intercept is 32 because this is the amount he starts with

jimthompson5910 (jim_thompson5910):

So John's equation is y = 1x + 32 which is the same as y = x + 32

jimthompson5910 (jim_thompson5910):

Because of your answer choices, you can easily see which one it is but I recommend going through the same steps to find William's equation

OpenStudy (madhatter):

ok thanks so much

jimthompson5910 (jim_thompson5910):

you're welcome